OpenStudy (anonymous):

A milling company wants to mix alfalfa, which contains 20% protein and wheat mids, which contains 15% protein to make cattle feed. The milling company wants to make 1000 pounds of cattle feed that contains 17% protein. How many pounds of alfalfa and how many pounds of wheat mids must be used?

OpenStudy (amistre64):

(amount) times (percentage) is the basic unit for this problem

OpenStudy (amistre64):

we have an amount of "a" at 20%. And we have an amount "w" at 15%. when we add these amount together they sould equal 1000 lbs total and have 17% protien.

OpenStudy (amistre64):

a + w = 1000; this means that: a = 1000-w a(.20) + w(.15) = 1000(.17) (1000-w)(.20) + .15w = 170 200 - .20w + .15w = 170 -.05w = -30 .05w = 30 w = 600 lbs a + 600 = 1000; so a = 400lbs lets see if it works..... 400(.20) + 600(.15) = 1000(.17) 80 + 90 = 170 170 = 170 ...its good :)